OpenStudy (anonymous):

Which of the following numbers has a value greater than 3? A. 31% B. 12/5 C. 3.0 D. 10/3

OpenStudy (anonymous):

lets see

OpenStudy (anonymous):

I thinks its either A or D

OpenStudy (anonymous):

31% = .31 so A is definitely not it

OpenStudy (anonymous):

It's D

OpenStudy (anonymous):

10/3 = 3.333...repeating

OpenStudy (anonymous):

So its D?

OpenStudy (anonymous):

12/5 = 2.4 so B isn't it

OpenStudy (anonymous):

Yes, it's d >_>

OpenStudy (anonymous):

3.0 = 3 so C isn't it

OpenStudy (anonymous):

So its D lol

OpenStudy (anonymous):

10/3 = 3.333333333 so it's D

OpenStudy (anonymous):

Ok thank you!

OpenStudy (anonymous):

10/3 > 3 by 1/3

OpenStudy (anonymous):

http://i.imgur.com/q3O8A.gif